### 20 Mar 2026 v2

# Tensor Train Decomposition-based Channel Estimation for MIMO-AFDM Systems with Fractional Delay and Doppler

Abstract
Affine Frequency Division Multiplexing (AFDM) has emerged as a promising chirp-based multicarrier technology for high-speed communication systems. By applying discrete affine Fourier transform (DAFT), AFDM has the potential to achieve significant improvements in spectral efficiency (SE) while maintaining a bit error rate (BER) performance comparable to that of OTFS. To fully exploit the diversity gain offered by AFDM, accurate channel estimation is essential. However, existing studies have mainly focused on the integer-delay-tap scenario and single-symbol pilot-based estimation. Since delay taps in practice are generally fractional, approximating them as integers not only degrades delay estimation accuracy but also severely affects Doppler frequency estimation. To address this problem, in this paper, we investigate channel estimation for multiple-input multiple-output (MIMO)-AFDM systems. A time-affine frequency (T-AF) domain pilot structure is proposed to exploit time-domain phase variations. By leveraging the rotational invariance property in the spatial and temporal domains, a channel estimation algorithm based on Vandermonde-structured tensor-train ($T$) decomposition is developed. The proposed algorithm demonstrates superior computational efficiency compared with state-of-the-art parameter estimation methods. Moreover, diverging from current studies, we derive the global Ziv-Zakai bound (ZZB) as an alternative parameter estimation error lower bound to the Cramér-Rao bound (CRB). Numerical results show that the derived ZZB provides tighter global performance characterization and successfully captures the threshold phenomenon in mean square error (MSE) performance in the low-SNR regime. Furthermore, the proposed algorithm achieves superior communication performance relative to the existing schemes, while offering a computational speedup, reducing the execution time by an order of magnitude compared to the state-of-the-art iterative algorithms.
Index Terms:
AFDM, tensor decomposition, channel estimation, embedded pilot, fractional delay and Doppler.
I Introduction
Ubiquitous connectivity is envisioned as one of the fundamental usage scenarios of future sixth-generation (6G) communication systems, with the objective of providing global connectivity and high-quality communication service. A critical challenge in achieving this vision lies in guaranteeing reliable communication in high-mobility scenarios, including vehicle-to-everything (V2X), unmanned aerial vehicles (UAVs), high-speed trains, and non-terrestrial networks (e.g., low Earth orbit (LEO) satellites). These applications generally involve extremely high relative velocities, which induce severe Doppler spreads. The existing multicarrier technologies, such as orthogonal frequency division multiplexing (OFDM), suffer from significant inter-carrier interference (ICI) in these scenarios, resulting in severe degradation of communication quality [1, 2, 3].
To fundamentally overcome the performance degradation of OFDM in high-mobility scenarios, researchers have turned to novel modulation waveforms. One of the promising approaches is the delay-Doppler ($D$) domain modulation, with orthogonal time frequency space (OTFS) being a representative technique [4, 5]. OTFS is a two-dimensional (2D) modulation scheme that transforms the multipath doubly selective channel into sparse and approximately time-invariant impulse responses in the delay-Doppler ($D$) domain, thereby improving communication reliability. However, OTFS also suffers from notable drawbacks. Firstly, OTFS requires block-wise modulation over multiple symbols, which introduces inherent latency into the system [6]. Secondly, due to its 2D modulation structure, OTFS symbols exhibit intrinsic 2D coupling, preventing symbol-by-symbol processing as in OFDM. Instead, the receiver must perform joint detection over 2D signal blocks, which significantly increases detection complexity [7].
Chirp signal-based waveform design has also attracted considerable attention. Recently, a novel waveform termed affine frequency division multiplexing (AFDM) has been proposed for next-generation wireless communications to mitigate performance degradation in high-mobility scenarios [8]. AFDM employs the inverse discrete affine Fourier transform (IDAFT) to map data symbols onto multiple orthogonal chirp signals. By properly configuring the affine parameters according to the maximum Doppler frequency, AFDM converts the doubly selective channel into a quasi-static and sparse representation in the one-dimensional discrete affine frequency domain, similar pirit to OTFS. It has been shown in [8] that AFDM can achieve optimal multipath diversity. Results reported in [8, 9] demonstrated that AFDM achieves bit error rate (BER) performance comparable to OTFS while significantly outperforming OFDM. Furthermore, different from the 2D modulation of OTFS, which needs to buffer multiple symbols for block processing. AFDM involves no inter-symbol coupling. Consequently, it is compatible with the current 5G new radio (NR) frame structure [10].
Nevertheless, the results reported in [8, 9] rely on the ideal assumption of perfect channel state information (CSI). In practical deployments, however, CSI is generally imperfect, which necessitates performance evaluation under practical channel estimation. In [11], a linear minimum mean square error (LMMSE) estimator was derived for single-input-single-output (SISO)-AFDM systems. In [12], an embedded pilot-aided diagonal reconstructability algorithm was proposed, which detects partial channel responses and reconstructs the entire AFDM channel by leveraging its diagonal reconstruction property. Nevertheless, the above methods did not explicitly estimate the underlying physical channel parameters, including the delay, Doppler and channel gain. In [13], the authors proposed a MIMO-AFDM channel estimation algorithm based on the embedded pilot structure, where the delay and Doppler of each propagation path were assumed to lie on the discrete grid. The authors of [14] proposed a superimposed pilot structure and an orthogonal matching pursuit (OMP)-based channel estimation algorithm. To address fractional Doppler estimation, [15] developed a sparse Bayesian learning (SBL)-based off-grid channel estimation algorithm, where hyperparameters are updated via the expectation maximization (EM) algorithm, resulting in high computational complexity. Furthermore, [16] proposed a grid-evolution SBL algorithm for integer-delay and fractional-Doppler estimation using the first-order Taylor expansion. However, the existing literature did not account for fractional delays. It is noteworthy that, within the AFDM input-output framework, the discrete affine frequency domain peak of a propagation path is determined by a linear combination of its delay and Doppler frequency. With a sufficiently large number of subcarriers, neglecting the fractional component of delay introduces negligible delay estimation error yet leads to considerable Doppler frequency estimation deviation. While such deviation can be compensated via simple phase correction in single-path channels, it becomes destructive in multipath scenarios, as distinct Doppler frequencies across paths produce varying phase errors that cannot be jointly corrected with a single global factor [17].
Motivated by the aforementioned challenges, in this paper, we investigate the pilot structure design and channel estimation for MIMO-AFDM systems with the consideration of fractional delay and Doppler. The main contributions of this paper are summerized as follows:
  • We analyze the input-output relationship of MIMO-AFDM systems under fractional delay and Doppler conditions. Based on this analysis, we design an embedded pilot structure arranged along the time dimension, which differs fundamentally from the existing single-symbol pilot estimation schemes.
  • Based on the proposed pilot structure, we model the received signal of MIMO-AFDM and formulate the channel estimation problem. By exploiting the rotational invariance property of MIMO-AFDM signals in space and time domains, as well as the impulse response magnitude characteristics in the affine frequency domain, we propose an efficient channel estimation algorithm based on Vandermonde-structured tensor-train ($T$) decomposition. The proposed algorithm achieves significantly lower computational complexity than the existing iterative algorithms.
  • To further evaluate the performance of parameter estimation, we provide analysis of the lower bound on the mean squared error (MSE) for channel parameter estimation. Different from the existing studies that adopt the Cramér-Rao bound (CRB) as benchmark, we derive the Ziv-Zakai bound (ZZB) as an alternative reference for the MSE lower bound. Compared with the CRB, the ZZB offers superior global tightness and can effectively capture the threshold effect in parameter estimation under low signal-to-noise ratio (SNR) conditions.
  • Simulation results demonstrate that the proposed algorithm achieves superior parameter estimation performance in the medium-to-high SNR regime. In terms of communication performance metrics, including normalized mean squared error (NMSE), bit error rate (BER), and spectral efficiency (SE), the proposed scheme consistently outperforms benchmark methods. In addition, it achieves a one-to-two orders-of-magnitude reduction in computational complexity relative to the existing algorithms.
The rest of this paper is organized as follows. Section II describes the MIMO-AFDM system model. In Section III, we propose a T-AF domain embedded pilot structure and channel estimation algorithm. The ZZB analysis is provided in Section IV. Section V provides simulation results. Finally, conclusions are drawn in Section VI.
II MIMO-AFDM System Model
Figure 1: An example of MIMO-AFDM systems.
Consider a MIMO-AFDM system shown in Fig. 1. The mobile station (MS) is equipped with a single antenna, while the base station (BS) employs a uniform linear array (ULA) consisting of $N_{\mathrm{BS}}$ antennas. The total number of subcarriers is $M$. The sampling interval is given by $T_{s}=\frac{1}{M\Delta f}$, where $\Delta f$ represents the subcarrier spacing. The duration of the chirp-periodic prefix (CPP) is $T_{\mathrm{CPP}}=M_{\mathrm{CPP}}T_{s}$. Let $T_{\mathrm{sym}}$ and $T$ denote the symbol duration with and without the CPP, respectively. Consequently, the symbol duration without CPP is $T=\frac{1}{\Delta f}$ and $T=\frac{M}{M+M_{\mathrm{CPP}}}T_{\mathrm{sym}}$. Let $x_{n,m}$ denote the symbol transmitted on the $m$-th subcarrier of the $n$-th AFDM symbol. Denote $\mathbf{x}_{n}=\left[x_{n,0},x_{n,1},...,x_{n,M-1}\right]^{\mathrm{T}}$ and the corresponding modulated signal as $\mathbf{s}_{n}=\left[s_{n,0},s_{n,1},...,s_{n,M-1}\right]^{\mathrm{T}}$. According to [8], the modulated signal of the $n$-th AFDM symbol is given by
$\mathbf{s}_{n}=\mathbf{\Lambda}_{c1}^{\mathrm{H}}\mathbf{F}_{M}^{\mathrm{H}}\mathbf{\Lambda}_{c2}^{\mathrm{H}}\mathbf{x}_{n},$ (1)
where $\mathbf{\Lambda}_{c}=\mathrm{Diag}\left(e^{-j2\pi cm^{2}},m=0,1,...,M-1\right)$, and $\mathbf{F}_{M}$ is the $M\times M$ DFT matrix. The analog time domain symbols with rectangular waveform shaping can be then written as
$s\left(t\right)=\sum_{n}{}\sum_{m}{}x_{n,m}\phi_{m}\left(t-nT_{\mathrm{sym}}\right)g_{\mathrm{tx}}\left(t-nT_{\mathrm{sym}}\right),$ (2)
where the AFDM chirp basis function $\phi_{m}$ for the $m$-th subcarrier is given by
$\phi_{m}\left(t\right)$ $=e^{j\left(\frac{d}{2b}\left(t-T_{\mathrm{CPP}}\right)^{2}+2\pi\left(m\Delta f\right)\left(t-T_{\mathrm{CPP}}\right)+\frac{a}{2b}\left(m\Delta f\right)^{2}\right)},$ (3)
$\qquad\qquad\qquad\qquad\qquad\qquad 0\leq t\leq T_{\mathrm{sym}},$
where $\left(a,b,c,d\right)$ are the parameters of affine transform. The rectangular waveform function $g_{\mathrm{tx}}\left(t\right)$ are given by
[…]
Note that the CPP has already been included in (3). In AFDM, we have $c_{1}=\frac{d}{4\pi b}T_{s}^{2}$ and $c_{2}=\frac{a}{4\pi b}\Delta f^{2}$. Generally, by setting $c_{1}=\frac{2\left(\alpha_{\max}+k_{\nu}\right)+1}{2M}$ and $c_{2}$ to a value much smaller than $\frac{1}{2M}$ can AFDM achieve full diversity [8], where $\alpha_{\max}=\nu_{\max}T$ and $\nu_{\max}$ denotes the maximum Doppler frequency, and $k_{\nu}$ denotes the spacing factor. The received signal $\mathbf{r}\left(t\right)$ through the multipath time-varying channel at the BS is given by
$\mathbf{r}\left(t\right)$ $=\iiint{}\mathbf{a}_{N_{\mathrm{BS}}}\left(\theta\right)h\left(\theta,\tau,\nu\right)e^{j2\pi\nu\left(t-\tau\right)}$ (5)
$\quad\times s\left(t-\tau\right)\mathrm{d}\tau\mathrm{d}\nu\mathrm{d}\theta+\mathbf{n}\left(t\right),$
where the space-delay-Doppler channel response $h\left(\theta,\tau,\nu\right)$ and the array steering vector $\mathbf{a}_{N_{\mathrm{BS}}}\left(\theta\right)$ are defined as
$h\left(\theta,\tau,\nu\right)=\sum_{i=1}^{P}{\alpha_{i}}{\bm{\delta}}\left(\theta-\theta_{i}\right){\bm{\delta}}\left(\tau-\tau_{i}\right){\bm{\delta}}\left(\nu-\nu_{i}\right),$ (6)
where $P$ denotes the number of propagation paths and $\mathbf{a}_{X}\left(\theta\right)=\left[1,e^{j2\pi\frac{d}{\lambda}\cos\theta},...,e^{j2\pi\frac{d}{\lambda}\left(X-1\right)\cos\theta}\right]^{\mathrm{T}}$. $\mathbf{n}\left(t\right)$ denotes the additive white Gaussian noise vector and each element follows an i.i.d complex Gaussian distribution $\mathcal{C}\mathcal{N}\left(0,\sigma_{\mathrm{n}}^{2}\right)$. The $\alpha_{i}$ $\theta_{i}$, $\tau_{i}$ and $\nu_{i}$ represent the complex channel gain, the arrival of angle (AOA), delay, and the Doppler frequency of the $i$-th propagation path, respectively. The delay and Doppler taps for the $i$-th path can be denoted as
$\tau_{i}=\frac{l_{i}+\iota_{i}}{M\Delta f},\nu_{i}=\frac{k_{i}+\kappa_{i}}{T},$ (7)
where $l_{i}$ and $k_{i}$ are the integer indices of the delay tap and Doppler tap for the $i$-th propagation path, and $\iota_{i}\in(-\frac{1}{2},\frac{1}{2}]$ and $\kappa_{i}\in(-\frac{1}{2},\frac{1}{2}]$ are the fractional delay and Doppler for the $i$-th propagation path, respectively. After sampling, the received signals are demodulated via the discrete affine Fourier transform (DAFT) [8]. We have the following theorem to characterize the input-output relationship in the spatial-time-affine frequency (STAF) domain.
Theorem 1
The input-output relationship of the MIMO-AFDM system can be expressed as
$Y\left[n_{\mathrm{BS}},n,m\right]$ $=\sum_{m^{\prime}=0}^{M-1}{}x_{n,m^{\prime}}\sum_{i=1}^{P}{}\tilde{\alpha}_{i}\left[\mathbf{a}_{N_{\mathrm{BS}}}\left(\theta\right)\right]_{n_{\mathrm{BS}}}$ (8)
$\quad\times e^{j2\pi\nu_{i}nT_{\mathrm{sym}}}H_{i}\left[m,m^{\prime}\right]+N\left[n_{\mathrm{BS}},n,m\right],$
where $N\left[n_{\mathrm{BS}},n,m\right]$ is the additive noise and
$H_{i}\left[m,m^{\prime}\right]$ $=e^{j\frac{2\pi}{M}\left(Mc_{1}\left(l_{i}+\iota_{i}\right)^{2}-m^{\prime}\left(l_{i}+\iota_{i}\right)+Mc_{2}\left({m^{\prime}}^{2}-m^{2}\right)\right)}$
$\quad\times\frac{e^{-j2\pi\left(m-m^{\prime}-\mathrm{loc}_{i}\right)}-1}{Me^{-j\frac{2\pi}{M}\left(m-m^{\prime}-\mathrm{loc}_{i}\right)}-M}$
$\tilde{\alpha}_{i}$ $=\alpha_{i}e^{j2\pi\nu_{i}\left(\frac{M_{\mathrm{CP}}}{M}T-\tau_{i}\right)},$
$\mathrm{loc}_{i}$ $=\nu_{i}T-2Mc_{1}\left(l_{i}+\iota_{i}\right).$ (9)
respectively.
Proof: The proof is provided in Appendix A. $\blacksquare$
From the input-output relationship (8), the following three conclusions can be drawn.
[…]
2. The AFDM effective channel exhibits diagonal reconstructability [12]. Specifically, the complete affine frequency domain channel can be reconstructed once the channel parameters are estimated from the support set associated with a certain symbol on the $m^{\prime}$ subcarrier.
3. When the integer delay constraint is relaxed, the number of unknown parameters to be estimated increases. Consequently, the maximum likelihood estimation algorithm for the delay and Doppler using a single AFDM symbol becomes imprecise [8, 18, 19].
The above analysis suggests that single-symbol estimation is insufficient for resolving fractional delay and Doppler shifts accurately. The assumption of integer delay neglects the fractional part $\left\{\iota_{i}\right\}$, which not only results in a coarse delay estimation but also leads to severe deviations in the Doppler frequency estimation $\left|\Delta\nu_{i}\right|=\left|2Mc_{1}\iota_{i}/T\right|$. Typically, $c_{1}=\frac{2\left(\alpha_{\max}+k_{\nu}\right)+1}{2M}$, and this error is substantial and cannot be neglected. Furthermore, from a positioning perspective, this assumption can result in decimeter-level positioning errors. For example, in the context of 5G NR numerologies, the subcarrier spacing is generally set to $2^{\mu}\cdot 15$ kHz, where $\mu=0,1,...,6$. With 1024 subcarriers and a 30 kHz spacing, the system achieves a fine delay resolution of 32.55 ns, equivalent to a spatial resolution of 9.76 m. Thus, for both communication and positioning tasks, it is imperative to incorporate fractional delay and Doppler into the system model.
III Pilot Structure Design And Channel Estimation Algorithm
In this section, we first introduce the proposed MIMO-AFDM embedded pilot structure and formulate the channel estimation problem. Then, by exploiting the multidimensional structure and inherent low-rank properties of the MIMO-AFDM signals, we develop a computationally efficient channel estimation framework based on Vandermonde-structured tensor-train ($T$) decomposition tailored to the proposed pilot structure.
III-A Proposed T-AF Embedded Pilot Structure
Figure 2: Proposed time-affine frequency domain embedded pilot structure.
[…]
Let $m_{\mathrm{pilot}}$ and NN denote the position of the embedded pilot symbol in each AFDM symbols and the number of AFDM symbols that are inserted pilot symbols, respectively. By collecting all the pilot signals received on the $n_{\mathrm{BS}}$-th antenna, we have
$\widehat{\mathbf{Y}}_{n_{\mathrm{BS}}}=\sum_{i=0}^{P-1}{}\tilde{\alpha}_{i}e^{j2\pi\frac{d}{\lambda}\left(n_{\mathrm{BS}}-1\right)\cos\theta_{i}}\mathbf{b}_{i}\mathbf{c}_{i}^{\mathrm{T}}+\mathbf{N}_{n_{\mathrm{BS}}},$ (10)
where
$\mathbf{b}_{i}\left(n\right)$ $=e^{j2\pi\nu_{i}nT_{\mathrm{sym}}}$ (11)
$\mathbf{c}_{i}\left(m\right)$ $=H_{i}\left[m,m_{\mathrm{pilot}}\right]X\left[m_{\mathrm{pilot}}\right].$ (12)
By further stacking the signals in the space-time-affine frequency domain, we can obtain the following tensor model in the CPD format as
$\widehat{\bm{\mathcal{Y}}}=\bm{\mathcal{S}}\times_{1}\mathbf{A}_{1}\times_{2}\mathbf{A}_{2}\times_{3}\mathbf{A}_{3}+\bm{\mathcal{N}},$ (13)
where the core tensor $\bm{\mathcal{S}}$ and the factor matrices $\mathbf{A}_{i},i=1,2,3$ are given by
$\bm{\mathcal{S}}$ $=\bm{\mathcal{D}}\left(\left[\tilde{\alpha}_{1},\tilde{\alpha}_{2},...,\tilde{\alpha}_{P}\right]^{\mathrm{T}}\right)\in\mathbb{C}^{P\times P\times P}$
$\mathbf{A}_{1}$ $=\left[\mathbf{a}_{N_{\mathrm{BS}}}\left(\theta_{1}\right),\mathbf{a}_{N_{\mathrm{BS}}}\left(\theta_{2}\right),...,\mathbf{a}_{N_{\mathrm{BS}}}\left(\theta_{P}\right)\right]\in\mathbb{C}^{N_{\mathrm{BS}}\times P}$
$\mathbf{A}_{2}$ $=\left[\mathbf{b}_{1},\mathbf{b}_{2},...,\mathbf{b}_{P}\right]\in\mathbb{C}^{N\times P}$
$\mathbf{A}_{3}$ $=\left[\mathbf{c}_{1},\mathbf{c}_{2},...,\mathbf{c}_{P}\right]\in\mathbb{C}^{M_{\mathrm{region}}\times P},$ (14)
where the entry in $\mathbf{b}_{i}$ and $\mathbf{c}_{i}$ are given by $\mathbf{b}_{i}\left(n\right)=e^{j2\pi\nu_{i}nT_{\mathrm{sym}}}$ and $\mathbf{c}_{i}\left(m\right)=H_{i}\left[m,m_{\mathrm{pilot}}\right]X\left[m_{\mathrm{pilot}}\right]$, respectively. From (III-A), it can be seen that the three dimensions of the tensor $\widehat{\bm{\mathcal{Y}}}$ contain the angle information $\left\{\theta_{i}\right\}$, the Doppler information $\left\{\nu_{i}\right\}$, and the intercept $\left\{\mathrm{loc}_{i}\right\}$, respectively. Therefore, MIMO-AFDM channel estimation under fractional delay and Doppler can be transformed into the tensor decomposition and parameter estimation problem.
III-B Proposed Tensor Train based Channel Estimation Method
In this subsection, we focus on the decomposition of $\widehat{\bm{\mathcal{Y}}}$ and channel parameters estimation.
III-B1 Factor Matrices Retrieval
From (13) and (III-A), it can be seen that the tensor decomposition problem in (13) can be solved by canonical polyadic decomposition (CPD). Representative methods to solve such problem include a series of alternating iterative algorithms derived from the core ALS approach, such as the rectified ALS (RecALS) [20], the SVD-employed ALS [21] and tensorial minimum mean square error (T-MMSE) [22]. However, the robustness and convergence of ALS face significant challenges when the tensor rank exceeds 2 [23], and the computational efficiency of ALS based algorithms remains insufficient for high-mobility scenarios. As a novel tensor decomposition method, $T$-decomposition [24] decomposes a high-dimensional tensor into multiple three-dimensional tensors connected in a chain structure, which offers greater flexibility and higher computational efficiency compared to CPD. Moreover, the decomposition error of the $T$-decomposition can be strictly upper bounded by the sum of squared singular values discarded in the SVD of each dimension, effectively addressing the insufficient convergence and robustness of ALS. Without loss of generality, in the following derivations, we use $\bm{\mathcal{Y}}=\bm{\mathcal{S}}\times_{1}\mathbf{A}_{1}\times_{2}\mathbf{A}_{2}\times_{3}\mathbf{A}_{3}$ to denote the noiseless signal model and $\widehat{\bm{\mathcal{Y}}}$ to denote the noisy signal model.
The $T$-decomposition represents tensor $\bm{\mathcal{Y}}$ as a chain-like product of three tensor cores as
$\bm{\mathcal{Y}}$ $=\sum_{i_{1}=1}^{P}{}\sum_{i_{2}=1}^{P}{}\mathbf{G}_{1}\left(:,i_{1}\right)\bm{\mathcal{G}}_{2}\left(i_{1},:,i_{2}\right)\mathbf{G}_{3}\left(i_{2},:\right)$
$=\mathbf{G}_{1}\times_{2}^{1}\bm{\mathcal{G}}_{2}\times_{3}^{2}\mathbf{G}_{3},$ (15)
where $\mathbf{G}_{1}\in\mathbb{C}^{N_{\mathrm{BS}}\times P}$, $\bm{\mathcal{G}}_{2}\in\mathbb{C}^{P\times N\times P}$ and $\mathbf{G}_{3}\in\mathbb{C}^{P\times M_{\mathrm{region}}}$ denote the head matrix, the core tensor and the tail matrix, respectively. In this paper, the $T$-SVD is applied to reconstruct the $T$-cores $\mathbf{G}_{1}$, $\bm{\mathcal{G}}_{2}$ and $\mathbf{G}_{3}$ [25], which can be represented by
$\mathbf{U}_{1}\mathbf{V}_{1}$ $=\mathbf{A}_{1}\left(\mathbf{A}_{3}\odot\mathbf{A}_{2}\right)^{\mathrm{T}}=\mathbf{U}_{1}\mathbf{M}_{1}\left(\mathbf{A}_{3}\odot\mathbf{A}_{2}\right)^{\mathrm{T}}$
$\mathbf{G}_{1}$ $=\mathbf{A}_{1}\mathbf{M}_{1}^{-1}$
$\mathbf{V}_{1}$ $=\mathbf{M}_{1}\left(\mathbf{A}_{3}\odot\mathbf{A}_{2}\right)^{\mathrm{T}}$ (16)
$\mathbf{U}_{2}\mathbf{V}_{2}$ $=\mathrm{reshape}\left(\mathbf{V}_{1};PN,M_{\mathrm{region}}\right)$
$=\left(\mathbf{A}_{2}\odot\mathbf{M}_{1}\right)\mathbf{A}_{3}^{\mathrm{T}}=\mathbf{U}_{2}\mathbf{M}_{2}\mathbf{A}_{3}^{\mathrm{T}}$
$\mathbf{U}_{2}$ $=\left(\mathbf{A}_{2}\odot\mathbf{M}_{1}\right)\mathbf{M}_{2}^{-1},$
[…]
$\mathbf{G}_{3}$ $=\mathbf{V}_{2}=\mathbf{M}_{2}\mathbf{A}_{3}^{\mathrm{T}},$ (18)
where $\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ are the non-singular change-of-bias matrices. The singular value matrix is assumed to be incorporated into the right singular matrix. It is observed from the derivation that the CPP factor matrices are intrinsically embedded within the subspaces of the $T$-cores. Specifically, the column spaces of $\mathbf{A}_{1}$, $\mathbf{A}_{2}$ and $\mathbf{A}_{3}$ coincides with the column space of $\mathbf{G}_{1}$, the subspace spanned by the mode-2 fibers of $\bm{\mathcal{G}}_{2}$ and the row space of $\mathbf{G}_{3}$, respectively. Given the Vandermonde nature of $\mathbf{A}_{1}$ and $\mathbf{A}_{3}$, their defining parameters can be retrieved from the subspaces of $\mathbf{G}_{1}$ and the subspaces of the mode-2 unfolding of $\mathbf{G}_{2}$ via ESPRIT. Nevertheless, this approach incurs a column ambiguity. Since the parameters are estimated separately, there is no inherent mechanism to align the path indices, potentially resulting in mismatched pairs of channel parameters for the propagation paths.
To resolve the ambiguity in parameter pairing, we leverage the intrinsic properties of the Vandermonde matrices to perform a sequential reconstruction of three factor matrices. According to (16), we have
$\underline{\mathbf{G}}_{1}\mathbf{M}_{1}$ $=\underline{\mathbf{A}}_{1}$ (19)
$\overline{\mathbf{G}}_{1}\mathbf{M}_{1}$ $=\overline{\mathbf{A}}_{1}=\underline{\mathbf{G}}_{1}\mathbf{M}_{1}\mathbf{Z}_{1},$
where $\underline{\mathbf{G}}_{1}$ and $\overline{\mathbf{G}}_{1}$ represent the matrix ${\mathbf{G}}_{1}$ with its last row removed and its first row removed, respectively. Then the generators $\left\{z_{1,r}\right\}$ of Vandermonde matrix $\mathbf{A}_{1}$ can be estimated via the ESPRIT algorithm
$\underline{\mathbf{G}}_{1}^{\dagger}\overline{\mathbf{G}}_{1}=\widehat{\mathbf{M}}_{1}\mathbf{Z}_{1}\widehat{\mathbf{M}}_{1}^{-1},$ (20)
where $\widehat{z}_{1,r}=\mathbf{Z}_{1}\left(r,r\right)/\left|\mathbf{Z}_{1}\left(r,r\right)\right|$. Consequently, the Vandermonde factor matrix $\mathbf{A}_{1}$ can be reconstructed with permutation ambiguity as $\widehat{\mathbf{A}}_{1}=\mathbf{A}_{1}\mathbf{\Pi}$, where $\mathbf{\Pi}$ is a permutation matrix. Next, we use the estimated $\widehat{\mathbf{A}}_{1}$ to construct a pseudoinverse matrix and then recover $\mathbf{A}_{2}$. Multiplying $\mathbf{G}_{1}$ and the mode-2 unfolding of $\bm{\mathcal{G}}_{2}$, we have
$\mathbf{G}_{1}\left[\bm{\mathcal{G}}_{2}\right]_{1}=\mathbf{A}_{1}\left(\mathbf{M}_{2}^{-\mathrm{T}}\odot\mathbf{A}_{2}\right)^{\mathrm{T}}.$ (21)
Left-multiplying the term by the pseudoinverse of $\widehat{\mathbf{A}}_{1}$ yields
$\widehat{\mathbf{A}}_{1}^{\dagger}\mathbf{G}_{1}\left[\bm{\mathcal{G}}_{2}\right]_{1}$ $=\mathbf{\Pi}^{-1}\left(\mathbf{M}_{2}^{-\mathrm{T}}\odot\mathbf{A}_{2}\right)^{\mathrm{T}}$
$=\left(\mathbf{M}_{2}^{-\mathrm{T}}\mathbf{\Pi}\odot\mathbf{A}_{2}\mathbf{\Pi}\right)^{\mathrm{T}}$
$=\left(\widehat{\mathbf{M}}_{2}^{-\mathrm{T}}\odot\widehat{\mathbf{A}}_{2}\right)^{\mathrm{T}}.$ (22)
Each row in $\left(\mathbf{M}_{2}^{-\mathrm{T}}\mathbf{\Pi}\odot\mathbf{A}_{2}\mathbf{\Pi}\right)^{\mathrm{T}}$ is a vector Kronecker product $\left(\widehat{\mathbf{M}}_{2}^{-\mathrm{T}}\odot\widehat{\mathbf{A}}_{2}\right)_{r,:}^{\mathrm{T}}=\left(\widehat{\mathbf{M}}_{2}^{-\mathrm{T}}\right)_{:,r}^{\mathrm{T}}\otimes\left(\widehat{\mathbf{A}}_{2}\right)_{:,r}^{\mathrm{T}}$. By reshapeing each row of the Khatri-Rao product $\left(\widehat{\mathbf{M}}_{2}^{-\mathrm{T}}\right)_{:,r}^{\mathrm{T}}\otimes\left(\widehat{\mathbf{A}}_{2}\right)_{:,r}^{\mathrm{T}}$ into matrix ${\mathbf{a}}_{2,r}\left(\mathbf{M}_{2}^{-\mathrm{T}}\right)_{:,r}^{\mathrm{T}}$, the factor matrix $\mathbf{A}_{2}$ and the nonsingular change-of-basis matrix $\mathbf{M}_{2}$ can be recovered by solving the following problem
$\left\{\widehat{\mathbf{a}}_{2,r},\left(\widehat{\mathbf{M}}_{2}^{-\mathrm{T}}\right)_{:,r}^{\mathrm{T}}\right\}$
$\qquad=\mathop{\mathrm{arg}\min}\limits_{\widehat{\mathbf{a}}_{2,r},\left(\widehat{\mathbf{M}}_{2}^{-\mathrm{T}}\right)_{:,r}^{\mathrm{T}}}\left\|\widehat{\mathbf{a}}_{2,r}\left(\mathbf{M}_{2}^{-\mathrm{T}}\right)_{:,r}^{\mathrm{T}}-\widehat{\mathbf{a}}_{2,r}\left(\widehat{\mathbf{M}}_{2}^{-\mathrm{T}}\right)_{:,r}^{\mathrm{T}}\right\|.$ (23)
The solution can be obtained by performing SVD of $\widehat{\mathbf{a}}_{2,r}\left(\mathbf{M}_{2}^{-\mathrm{T}}\right)_{:,r}^{\mathrm{T}}$. The estimated $\widehat{\mathbf{A}}_{2}$ and $\widehat{\mathbf{M}}_{2}^{-\mathrm{T}}$ exhibit permutation and scaling ambiguity as
$\widehat{\mathbf{A}}_{2}$ $=\mathbf{A}_{2}\mathbf{\Pi\Delta}^{-1},$ (24)
$\widehat{\mathbf{M}}_{2}^{-\mathrm{T}}$ $=\mathbf{M}_{2}^{-\mathrm{T}}\mathbf{\Pi\Delta},$ (25)
where $\bm{{\bm{\delta}}}$ is a diagnal scaling ambiguity matrix. Finally, the factor matrix ${\mathbf{A}}_{3}$ with ambiguity can be obtained as
$\widehat{\mathbf{A}}_{3}=\mathbf{G}_{3}^{\mathrm{T}}\widehat{\mathbf{M}}_{2}^{-\mathrm{T}}=\mathbf{A}_{3}\mathbf{\Pi\Delta}.$ (26)
III-B2 Channel Parameters Estimation
The channel parameters are estimated from the recovered factor matrices. The angles of arrival (AoAs) and Doppler frequencies can be directly estimated by exploiting the rotational invariance property of the factor matrices.
$\widehat{\theta}_{r}$ $=\mathrm{a}\cos\left(\frac{\lambda}{2\pi d}\angle\left(\mathbf{Z}_{1}\right)_{r,r}\right),$ (27)
$\widehat{\nu}_{r}$ $=\frac{1}{2\pi T_{\mathrm{sym}}}\angle\left(\widehat{\mathbf{A}}_{2}^{\dagger}\left(1:N-1,r\right)\widehat{\mathbf{A}}_{2}\left(2:N,r\right)\right),$ (28)
respectively.
We then focus on the estimation of delays. According to (1), for $m^{\prime}=m_{\mathrm{pilot}}$ and assuming zero delay and Doppler frequency, the peak occurs at $m=m^{\prime}=m_{\mathrm{pilot}}$. As the delay increases, the main lobe index $m$ decreases relative to $m^{\prime}=m_{\mathrm{pilot}}$. Let $m_{i}$ denote the fractional index corresponding to the unique peak observed in the affine frequency domain, where
$m_{r}$ $=m_{\mathrm{pilot}}+\nu_{r}T-2Mc_{1}\left(l_{r}+\iota_{r}\right)$ (29)
$=\mathrm{round}\left(m_{r}\right)+\tilde{m}_{r},$
with $\tilde{m}_{r}\in\left[-\frac{1}{2},\frac{1}{2}\right)$. Let $m_{r}^{\prime}$ and $m_{r}^{\prime\prime}$ be the indices of the main lobe and the largest side lobe in the affine frequency domain along the $r$-th propagation path, we have
$m_{r}^{\prime}$ $=\mathrm{round}\left(m_{r}\right)$ (30)
[…]
Then, according to (1), we have
$\frac{\left|H_{i}\left[m_{r}^{\prime},m_{\mathrm{pilot}}\right]\right|}{\left|H_{i}\left[m_{r}^{\prime\prime},m_{\mathrm{pilot}}\right]\right|}$ $=\frac{\left|\sin\left(-\frac{\pi}{M}\left(m_{r}^{\prime}-m_{r}^{\prime\prime}-\tilde{m}_{r}\right)\right)\right|}{\left|\sin\left(-\frac{\pi}{M}\tilde{m}_{r}\right)\right|}$
$\approx\frac{\left|m_{r}^{\prime}-m_{r}^{\prime\prime}-\tilde{m}_{r}\right|}{\left|\tilde{m}_{r}\right|}.$ (32)
The indices $m_{r}^{\prime}$ and $m_{r}^{\prime\prime}$ correspond to the row indices of the factor matrices $\widehat{\mathbf{A}}_{3}$ are given by
$m_{\mathrm{p}1,r}^{\prime}$ $=\left(m_{\mathrm{pilot}}-m_{\mathrm{region}}+1\right)+m_{r}^{\prime},$ (33)
$m_{\mathrm{p}2,r}^{\prime\prime}$ $=\left(m_{\mathrm{pilot}}-m_{\mathrm{region}}+1\right)+m_{r}^{\prime\prime}.$ (34)
Based on (III-B2), (33) and (34), the fractional part $\tilde{m}_{r}$ can be estimated as
$\widehat{\tilde{m}}_{r}\approx\frac{\left(m_{\mathrm{p}2,r}^{\prime\prime}-m_{\mathrm{p}1,r}^{\prime}\right)\left|\widehat{\mathbf{A}}_{3}\left(m_{\mathrm{p}2,r}^{\prime\prime},r\right)\right|}{\left|\widehat{\mathbf{A}}_{3}\left(m_{\mathrm{p}1,r}^{\prime},r\right)\right|+\left|\widehat{\mathbf{A}}_{3}\left(m_{\mathrm{p}2,r}^{\prime\prime},r\right)\right|}.$ (35)
The fractional delay can be then calculated by
$\widehat{l}_{r}+\widehat{\iota}_{r}$ $=\frac{m_{\mathrm{pilot}}+\widehat{\nu}_{r}T-\left(m_{\mathrm{p}1,r}^{\prime}+\widehat{\tilde{m}}_{r}\right)}{2Mc_{1}}$ (36)
$=\frac{m_{\mathrm{region}}+\widehat{\nu}_{r}T-m_{\mathrm{p}1,r}-\widehat{\tilde{m}}_{r}-1}{2Mc_{1}}.$
Algorithm 1 Proposed Channel Estimation Algorithm
[…]
1: Perform $T$-SVD by (16), (17) and (18) to obtain the tensor cores $\mathbf{G}_{1}$, $\bm{\mathcal{G}}_{2}$ and $\mathbf{G}_{3}$.
2: Compute the EVD (20) and extract the generators $\widehat{z}_{1,r}=\mathbf{Z}_{1}\left(r,r\right)/\left|\mathbf{Z}_{1}\left(r,r\right)\right|$.
3: Recover the factor matrix $\widehat{\mathbf{A}}_{1}$ as $\widehat{\mathbf{a}}_{1,r}=\left[1,\widehat{z}_{1,r},...,\widehat{z}_{1,r}^{N_{\mathrm{BS}}-1}\right]^{\mathrm{T}}$.
4: Compute the multiplication (21) and (III-B1), and recover $\widehat{\mathbf{A}}_{2}$ and $\widehat{\mathbf{M}}_{2}^{-\mathrm{T}}$ by (23).
5: Recover $\widehat{\mathbf{A}}_{3}$ by (26) .
6: Estimate the AoAs $\left\{\widehat{\theta}_{r}\right\}$ and Doppler frequencies $\left\{\widehat{\nu}_{r}\right\}$ by (27) and (28), respectively.
7: Estimate the delay taps $\left\{\widehat{l}_{r}+\widehat{\iota}_{r}\right\}$ by (III-B2)-(36).
8: Compute the multiplications by (37) and estimate the channel gain $\left\{\widehat{\tilde{\alpha}}_{r}\right\}$ by (38).
[…]
Finally, we examine the estimation of channel gains. Note that $\widehat{\mathbf{A}}_{3}$ is subject to an ambiguity factor $\bm{{\bm{\delta}}}$, which should be eliminated first. Based on the previously estimated AoA and Doppler parameters, we firstly reconstruct the factor matrices $\mathbf{A}_{1}\mathbf{\Pi}$ and $\mathbf{A}_{2}\mathbf{\Pi}$ with column ambiguities. Then, defining $\bm{\mathcal{Z}}=\mathbf{G}_{0}\times_{2}^{1}\bm{\mathcal{G}}_{1}\times_{3}^{2}\mathbf{G}_{2}$, we have the $\mathbf{A}_{3}\mathbf{\Pi}$ without scaling ambiguity as follows:
$\left[\bm{\mathcal{Z}}\right]_{3}$ $=\mathbf{A}_{3}\left(\mathbf{A}_{2}\odot\mathbf{A}_{1}\right)^{\mathrm{T}}$ (37)
$\mathbf{A}_{3}\mathbf{S\Pi}$ $=\left[\bm{\mathcal{Z}}\right]_{3}\left(\left(\mathbf{A}_{2}\mathbf{\Pi}\odot\mathbf{A}_{1}\mathbf{\Pi}\right)^{\dagger}\right)^{\mathrm{T}},$
where $\mathbf{S}=\mathrm{Diag}\left(\left[\tilde{\alpha}_{1},\tilde{\alpha}_{2},...,\tilde{\alpha}_{P}\right]^{\mathrm{T}}\right)$. The channel gain can be estimated via (38), which is at the bottom of the next page. Compared to the method in [26], which estimates parameters to recover all factor matrices and then obtains the channel gains by computing the pseudo-inverse of their Khatri-Rao product, the approach proposed in (38) utilizes the main lobe of the affine frequency domain response of the $r$-th propagation path to estimate the channel gains, offering greater robustness against noise. The proposed algorithm is detailed in Algorithm 1.
III-C Complexity Analysis
In this subsection, the computational of the proposed algorithm is analyzed. The main complexity of the proposed algorithm lies in SVD and pseudoinverse of matrices. Specifically, the $T$-SVD process takes the complexity of $\mathcal{O}\left(N_{\mathrm{BS}}^{2}NM_{\mathrm{region}}+P^{2}N^{2}M_{\mathrm{region}}\right)$. The EVD of (20) and the pseudoinverse and multiplication (III-B1) are on the order of the complexity of $\mathcal{O}\left(N_{\mathrm{BS}}^{3}\right)$ and $\mathcal{O}\left(N_{\mathrm{BS}}^{2}P+N_{\mathrm{BS}}P^{2}N\right)$, respectively. Finally, the total complexity of the proposed algorithm is of order $\mathcal{O}\left(N_{\mathrm{BS}}^{2}NM_{\mathrm{region}}+P^{2}N^{2}M_{\mathrm{region}}+N_{\mathrm{BS}}^{2}P+N_{\mathrm{BS}}P^{2}N\right)$.
$\widehat{\tilde{\alpha}}_{r}=\frac{\left(\mathbf{A}_{3}\mathbf{S\Pi}\right)_{m_{\mathrm{p}1,r}^{\prime},r}}{e^{j\frac{2\pi}{M}\left(Mc_{1}\left(\widehat{l}_{r}+\widehat{\iota}_{r}\right)^{2}-m_{\mathrm{pilot}}\left(\widehat{l}_{r}+\widehat{\iota}_{r}\right)+Mc_{2}\left(m_{\mathrm{pilot}}^{2}-\left(m_{\mathrm{p}1,r}^{\prime},r\right)^{2}\right)\right)}\frac{e^{-j2\pi\left(m_{\mathrm{p}1,r}^{\prime}-m_{\mathrm{pilot}}-\widehat{\nu}_{r}T+2Mc_{1}\left(\widehat{l}_{r}+\widehat{\iota}_{r}\right)\right)}-1}{Me^{-j\frac{2\pi}{M}\left(m_{\mathrm{p}1,r}^{\prime}-m_{\mathrm{pilot}}-\widehat{\nu}_{r}T+2Mc_{1}\left(\widehat{l}_{r}+\widehat{\iota}_{r}\right)\right)}-M}}$ (38)
IV ZZB Anaysis
The lower bound of the MSE is essential for evaluating channel parameter estimation performance. Most existing researches [27, 28] utilized the CRB as the performance benchmark. Nevertheless, at low signal-to-noise ratios (SNRs), the CRB fails to provide a tight bound on the actual estimation performance. As a local non-Bayesian bound, it exhibits asymptotic tightness only under small estimation error conditions (high SNR scenarios). In numerical simulations, the estimation performance generally suffers from severe degradation as the SNR drops in the low SNR region. This phenomenon is known as the threshold effect [29, 30, 31]. In this section, we investigate the close-form ZZB for estimation of channel parameters in multipath propagation, including AoAs, Doppler frequencies, and delay taps. Without loss of generality, we assume that the signals received over different propagation paths are incoherent, and the channel parameters associated with each path are independently and identically distributed (i.i.d.) while being mutually distinct.
Define $\eta_{i}=l_{i}+\iota_{i}$ and $\bm{p}=\left[\bm{\theta}^{\mathrm{T}},\bm{\nu}^{\mathrm{T}},\bm{\eta}^{\mathrm{T}}\right]^{\mathrm{T}}$. The vectorized form of the tensor $\bm{\mathcal{Y}}$ is denoted by $\mathbf{y}$, and is given by
[…]
$=\sum_{i=1}^{P}{}\tilde{\alpha}_{i}\bm{\phi}\left(\theta_{i},\nu_{i},\eta_{i}\right)+\mathbf{n}$
$=\mathbf{\Phi}\left(\bm{\theta},\bm{\nu},\bm{\eta}\right)\bm{\alpha}+\mathbf{n},$ (39)
where
$\bm{\phi}\left(\theta_{i},\nu_{i},\eta_{i}\right)$ $=\mathbf{a}_{N_{\mathrm{BS}}}\left(\theta_{i}\right)\otimes\mathbf{b}_{N}\left(\nu_{i}\right)\otimes\mathbf{c}_{M_{\mathrm{pilot}}}\left(\nu_{i},\eta_{i}\right)$
$\bm{\Phi}\left(\bm{p}\right)$ $=\left[\bm{\phi}\left(\theta_{1},\nu_{1},\eta_{1}\right),\bm{\phi}\left(\theta_{2},\nu_{2},\eta_{2}\right),...,\bm{\phi}\left(\theta_{P},\nu_{P},\eta_{P}\right)\right],$
and $\mathbf{\alpha}=\left[\tilde{\alpha}_{1},\tilde{\alpha}_{2},...,\tilde{\alpha}_{P}\right]^{\mathrm{T}}$ and $\mathbf{n}=\mathrm{vec}\left(\bm{\mathcal{N}}\right)$. Since DAFT is a unitary transform, each element of the noise tensor $\bm{\mathcal{N}}$ still follows an i.i.d complex Gaussian distribution $\mathcal{C}\mathcal{N}\left(0,\sigma_{\mathrm{n}}^{2}\right)$. Given the channel parameters $\bm{p}$, the covariance matrix $\mathbf{R}_{\mathbf{y}|\bm{p}}$ is written as
$\mathbf{R}_{\mathbf{y}|\bm{p}}$ $=\mathbb{E}\left\{\mathbf{yy}^{\mathrm{H}}\right\}$
$=\mathbf{\Phi}\left(\bm{p}\right)\mathbf{\Sigma\Phi}\left(\bm{p}\right)+\sigma_{n}^{2}\mathbf{I}_{N_{\mathrm{BS}}NM_{\mathrm{region}}},$ (40)
where $\mathbf{\Sigma}=\mathbb{E}\left\{\bm{\alpha\alpha}^{\mathrm{H}}\right\}=\mathrm{diag}\left(\left[\sigma_{1,\mathrm{PL}}^{2},\sigma_{2,\mathrm{PL}}^{2},...,\sigma_{P,\mathrm{PL}}^{2}\right]^{\mathrm{T}}\right)$. The error correlation matrix is defined as
$\mathbf{R}_{\mathbf{\epsilon}}=\mathbb{E}\left\{\mathbf{\epsilon\epsilon}^{\mathrm{T}}\right\}=\mathbb{E}\left\{\left(\hat{\bm{p}}-\bm{p}\right)\left(\hat{\bm{p}}-\bm{p}\right)^{\mathrm{T}}\right\}.$ (41)
Based on the above definitions, the extended ZZB is given by [32, 33]
$\bm{w}^{\mathrm{T}}\mathbf{R}_{\bm{\epsilon}}\bm{w}$ $\geq\frac{1}{2}\int_{0}^{\infty}{}\max_{\bm{{\bm{\delta}}}:\bm{w}^{\mathrm{T}}\bm{{\bm{\delta}}}=h}\left[\int_{\mathbb{R}^{3P}}{}\left(f_{\bm{p}}\left(\bm{{\bm{\varphi}}}\right)+f_{\bm{p}}\left(\bm{{\bm{\varphi}}}+\bm{{\bm{\delta}}}\right)\right)\right$.
$\quad\times\left.P_{\min}\left(\bm{{\bm{\varphi}}},\bm{{\bm{\varphi}}}+\bm{{\bm{\delta}}}\right)\mathrm{d}\bm{{\bm{\varphi}}}\right]h\mathrm{d}h,$ (42)
where the normalized weight vector $\bm{w}\in\mathbb{R}^{3P}$, satisfying $\left\|\bm{w}\right\|_{2}=1$. $f_{\bm{p}}\left(\bm{{\bm{\varphi}}}\right)$ and $f_{\bm{p}}\left(\bm{{\bm{\varphi}}}+\bm{{\bm{\delta}}}\right)$ are the a priori probability density function (pdf) of $\bm{p}$ given $\bm{p}=\bm{{\bm{\varphi}}}$ and $\bm{p}=\bm{{\bm{\varphi}}}+\bm{{\bm{\delta}}}$, respectively, and $P_{\min}\left(\bm{{\bm{\varphi}}},\bm{{\bm{\varphi}}}+\bm{{\bm{\delta}}}\right)$ represents the minimum probability of error associated with the following binary hypothesis testing problem:
$\mathcal{H}_{0}:\bm{p}=\bm{{\bm{\varphi}}};$ $\bm{y}\sim f\left(\bm{y}|\bm{p}=\bm{{\bm{\varphi}}}\right),$
$\mathcal{H}_{1}:\bm{p}=\bm{{\bm{\varphi}}}+\bm{{\bm{\delta}}};$ $\bm{y}\sim f\left(\bm{y}|\bm{p}=\bm{{\bm{\varphi}}}+\bm{{\bm{\delta}}}\right),$ (43)
where $f\left(\mathbf{y}|\bm{p}=\bm{{\bm{\varphi}}}\right)$ and $f\left(\mathbf{y}|\bm{p}=\bm{{\bm{\varphi}}}+\bm{{\bm{\delta}}}\right)$ are the conditional pdf of $\mathbf{y}$ given $\bm{p}=\bm{{\bm{\varphi}}}$ and $\bm{p}=\bm{{\bm{\varphi}}}+\bm{{\bm{\delta}}}$, respectively, and the prior probabilities of the two hypotheses are given by
$\mathrm{Pr}\left(\mathcal{H}_{0}\right)$ $=\frac{f_{\bm{p}}\left(\bm{{\bm{\varphi}}}\right)}{f_{\bm{p}}\left(\bm{{\bm{\varphi}}}+\bm{{\bm{\delta}}}\right)+f_{\bm{p}}\left(\bm{{\bm{\varphi}}}\right)},$
$\mathrm{Pr}\left(\mathcal{H}_{1}\right)$ $=1-\mathrm{Pr}\left(\mathcal{H}_{0}\right).$ (44)
Then, the MSE of the $k$-th set of channel parameters in $\bm{p}$} can be lower bounded by
$\mathrm{MSE}\left(\bm{p}_{k}\right)$ $=\bm{w}_{k}^{\mathrm{T}}\mathbf{R}_{\bm{\epsilon}}\bm{w}_{k}$ (45)
$\geq\frac{1}{2}\int_{0}^{\infty}{}\max_{\bm{{\bm{\delta}}}:\bm{w}_{k}^{\mathrm{T}}\bm{{\bm{\delta}}}=h}\left[\int_{\mathbb{R}^{3P}}{}\left(f_{\bm{p}}\left(\bm{{\bm{\varphi}}}\right)+f_{\bm{p}}\left(\bm{{\bm{\varphi}}}+\bm{{\bm{\delta}}}\right)\right)\right$.
$\quad\left.\times P_{\min}\left(\bm{{\bm{\varphi}}},\bm{{\bm{\varphi}}}+\bm{{\bm{\delta}}}\right)\mathrm{d}\bm{{\bm{\varphi}}}\right]h\mathrm{d}h,$
where
[…]
and
$\bm{w}_{k}=\frac{1}{\sqrt{P}}\left[\mathbf{0}_{\left(k-1\right)P}^{\mathrm{T}},\mathbf{1}_{P}^{\mathrm{T}},\mathbf{0}_{\left(3-k\right)P}^{\mathrm{T}}\right]^{\mathrm{T}}.$ (47)
However, the expression of $P_{\min}\left(\bm{{\bm{\varphi}}},\bm{{\bm{\varphi}}}+\bm{{\bm{\delta}}}\right)\mathrm{d}\bm{{\bm{\varphi}}}$ is generally complicated, and directly substituting it into (44) usually makes it difficult to obtain an explicit closed-form expression. Fortunately, a lower bound for $P_{\min}\left(\bm{{\bm{\varphi}}},\bm{{\bm{\varphi}}}+\bm{{\bm{\delta}}}\right)\mathrm{d}\bm{{\bm{\varphi}}}$ for observation $\bm{y}$ is given by [34]
$P_{\min}\left(\bm{{\bm{\varphi}}},\bm{{\bm{\varphi}}}+\bm{{\bm{\delta}}}\right)$
$\quad\geq P\left(\bm{{\bm{\delta}}}\right)$
$\quad=e^{\left[\mu\left(s;\bm{{\bm{\delta}}}\right)+\frac{1}{8}\frac{\partial^{2}\mu\left(s;\bm{{\bm{\delta}}}\right)}{\partial s^{2}}\right]}\left.\mathcal{Q}\left(\frac{1}{2}\sqrt{\frac{\partial^{2}\mu\left(s;\bm{{\bm{\delta}}}\right)}{\partial s^{2}}}\right)\right|_{s=\frac{1}{2}},$ (48)
where $\mathcal{Q}\left(x\right)=\int_{x}^{\infty}{}\frac{1}{\sqrt{2\pi}}e^{-\frac{v^{2}}{2}}\mathrm{d}v$ is the Q-function and
$\mu\left(s;\mathbf{{\bm{\delta}}}\right)$ $=\ln\int{}f\left(\mathbf{y}|\mathbf{{\bm{\varphi}}}+\mathbf{{\bm{\delta}}}\right)^{s}f\left(\mathbf{y}|\mathbf{{\bm{\varphi}}}\right)^{1-s}\mathrm{d}\mathbf{y}$
$=\ln\int{}\frac{1}{\pi^{N_{\mathrm{BS}}NM_{\mathrm{pilot}}}\left|\mathbf{R}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}+\mathbf{{\bm{\delta}}}}\right|^{s}\left|\mathbf{R}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}}\right|^{1-s}}$
$\quad\times\exp\left(-\mathbf{y}^{\mathrm{H}}\left(s\mathbf{R}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}+\mathbf{{\bm{\delta}}}}^{-1}+\left(1-s\right)\mathbf{R}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}}^{-1}\right)\mathbf{y}\right)\mathrm{d}\mathbf{y}$
$=-\ln\left|s\mathbf{R}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}+\mathbf{{\bm{\delta}}}}^{-1}+\left(1-s\right)\mathbf{R}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}}^{-1}\right|$
$\quad-s\ln\left|\mathbf{R}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}+\mathbf{{\bm{\delta}}}}\right|-\left(1-s\right)\ln\left|\mathbf{R}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}}\right|$ (49)
is the cumulant generating function (CGF) of the log-likelihood ratio $\ln\frac{f\left(\mathbf{y}|\mathbf{{\bm{\varphi}}}+\mathbf{{\bm{\delta}}}\right)}{f\left(\mathbf{y}|\mathbf{{\bm{\varphi}}}\right)}$. We present the following theorem that gives the closed-form expression for $P\left(\mathbf{{\bm{\delta}}}\right)$:
Theorem 2
Based on the CGF given in (49), we have
[…]
where $\left[\mathbf{J}\right]_{ij}=\mathrm{Tr}\left(\mathbf{R}^{-1}\frac{\partial\mathbf{R}}{\partial p_{i}}\mathbf{R}^{-1}\frac{\partial\mathbf{R}}{\partial p_{j}}\right)$ is the Fisher matrix and
$P_{\mathrm{NA}}=e^{\psi\left(\mathbf{\bm{\rho}}\right)}\mathcal{Q}\left(\sqrt{\frac{1}{2}\sum_{i=0}^{P-1}{}\left(\frac{2N_{\mathrm{BS}}N\rho_{i}}{2+N_{\mathrm{BS}}N\rho_{i}}\right)^{2}}\right),$ (51)
where
$\psi\left(\mathbf{\bm{\rho}}\right)=\sum_{i=0}^{P-1}\ln\left[\frac{4\left(1+N_{\mathrm{BS}}N\rho_{i}\right)}{\left(2+N_{\mathrm{BS}}N\rho_{i}\right)^{2}}\right]+\left(\frac{N_{\mathrm{BS}}N\rho_{i}}{2+N_{\mathrm{BS}}N\rho_{i}}\right)^{2},$ (52)
and the $i$-th element of $\mathbf{\bm{\rho}}$ is the SNR of the $i$-th propagation path, which is definded as $\rho_{i}=\frac{\sigma_{i,\mathrm{PL}}^{2}}{\sigma_{n}^{2}}$. The asymptotic region is defined as $\mathbf{\Delta}=\left\{\mathbf{{\bm{\delta}}}|\mathbf{{\bm{\delta}}}^{\mathrm{T}}\mathbf{J{\bm{\delta}}}\geq 2\sum_{i=1}^{P}{}\left(\frac{2N_{\mathrm{BS}}N\rho_{i}}{2+N_{\mathrm{BS}}N\rho_{i}}\right)^{2}\right\}$.
Proof: The proof is provided in Appendix B. $\blacksquare$
Based on Theorem 2, (IV) is lower bounded by
$\frac{1}{2}\int_{\mathbb{R}^{3P}}{}\left(f_{\bm{p}}\left(\mathbf{{\bm{\varphi}}}\right)+f_{\bm{p}}\left(\mathbf{{\bm{\varphi}}}+\mathbf{{\bm{\delta}}}\right)\right)P_{\min}\left(\mathbf{{\bm{\varphi}}},\mathbf{{\bm{\varphi}}}+\mathbf{{\bm{\delta}}}\right)\mathrm{d}\mathbf{{\bm{\varphi}}}$ (53)
$\quad\geq P\left(\mathbf{{\bm{\delta}}}\right)\prod_{k=1}^{3}{}\prod_{i=1}^{P}{}\frac{\left(\zeta_{k}-{\bm{\delta}}_{k,i}\right)}{\zeta_{k}},$
where $\zeta_{k}$ is the range of the uniform distribution for the $k$-th group of channel parameters in $\bm{p}$. Hence, (42) reduces to
$\bm{w}^{\mathrm{T}}\mathbf{Rw}\geq\max_{\mathbf{{\bm{\delta}}}:\bm{w}_{k}^{\mathrm{T}}\mathbf{{\bm{\delta}}}=h}P\left(\mathbf{{\bm{\delta}}}\right)\prod_{k=1}^{3}{}\prod_{i=1}^{P}{}\frac{\left(\zeta_{k}-{\bm{\delta}}_{k,i}\right)}{\zeta_{k}}.$ (54)
Substituting (50) into (54), we relax the maximization problem in (42) as [35]
$\max_{\mathbf{{\bm{\delta}}}:\bm{w}_{k}^{\mathrm{T}}\mathbf{{\bm{\delta}}}=h}\frac{P\left(\mathbf{{\bm{\delta}}}\right)}{\prod_{k=1}^{3}{}\zeta_{k}^{P}}\prod_{k=1}^{3}{}\prod_{i=1}^{P}{}\left(\zeta_{k}-{\bm{\delta}}_{k,i}\right)$ (55)
$\approx\max_{\mathbf{{\bm{\delta}}}:\bm{w}_{k}^{\mathrm{T}}\mathbf{{\bm{\delta}}}=h}\frac{P_{\mathrm{NA}}}{\prod_{k=1}^{3}{}\zeta_{k}^{P}}\prod_{k=1}^{3}{}\prod_{i=1}^{P}{\left(\zeta_{k}-{\bm{\delta}}_{k,i}\right)}$
$\quad+\max_{\mathbf{{\bm{\delta}}}\in\Delta:\bm{w}_{k}^{\mathrm{T}}\mathbf{{\bm{\delta}}}=h}P_{\mathrm{A}}\left(\mathbf{{\bm{\delta}}}\right)-P_{\mathrm{NA}}.$
The second term on the right-hand side of the approximate equality is due to the fact that $\mathbf{{\bm{\delta}}}\rightarrow\mathbf{0}$ in the asymptotic region. According to the inequality of arithmetic and geometric means, we have the closed-form for the first maximization term as
$\max_{\mathbf{{\bm{\delta}}}:\bm{w}_{k}^{\mathrm{T}}\mathbf{{\bm{\delta}}}=h}\frac{P\left(\mathbf{{\bm{\delta}}}\right)}{\prod_{k=1}^{3}{}\zeta_{k}^{P}}\prod_{k=1}^{3}{}\prod_{i=1}^{P}{}\left(\zeta_{k}-{\bm{\delta}}_{k,i}\right)$ (56)
$\quad=P_{\mathrm{NA}}\prod_{k=1}^{3}{}\left(1-\frac{h}{\sqrt{P}\zeta_{k}}\right)^{P}.$
And the second maximization term is given by
$\max_{\mathbf{{\bm{\delta}}}\in\Delta:\bm{w}_{k}^{\mathrm{T}}\mathbf{{\bm{\delta}}}=h}P_{\mathrm{A}}\left(\mathbf{{\bm{\delta}}}\right)=\mathcal{Q}\left(\frac{1}{2\sqrt{\bm{w}_{k}^{\mathrm{T}}\mathbf{J}^{-1}\bm{w}_{k}}}h\right),$ (57)
where the maximum value is achieved at $\mathbf{{\bm{\delta}}}=\frac{\mathbf{J}^{-1}\bm{w}_{k}}{\bm{w}_{k}^{\mathrm{T}}\mathbf{J}^{-1}\bm{w}_{k}}h$. Substituting (55)-(51) into (54), we obtain an approximate closed-form expression for ZZB as
$\bm{w}_{k}^{\mathrm{T}}\mathbf{R}_{\mathbf{\epsilon}}\bm{w}_{k}$ $\geq P_{\mathrm{NA}}\int_{0}^{\sqrt{K}\zeta}{}\left(1-\frac{h}{\sqrt{P}\zeta_{k}}\right)^{P}h\mathrm{d}h$
$\quad+\int_{0}^{\tilde{h}}{}\left[\mathcal{Q}\left(\frac{1}{2\sqrt{\bm{w}_{k}^{\mathrm{T}}\mathbf{J}^{-1}\bm{w}_{k}}}h\right)-P_{\mathrm{NA}}\right]h\mathrm{d}h$
$\approx P_{\mathrm{NA}}\frac{12\bm{w}_{k}^{\mathrm{T}}\mathbf{R}_{\bm{p}}\bm{w}_{k}}{\left(K+1\right)\left(K+2\right)}$
$\quad+\int_{0}^{\tilde{h}}{}\frac{1}{\sqrt{2\pi}}e^{-\frac{h^{2}}{8\bm{w}_{k}^{\mathrm{T}}\mathbf{J}^{-1}\bm{w}_{k}}}\frac{h^{2}}{4\sqrt{\bm{w}_{k}^{\mathrm{T}}\mathbf{J}^{-1}\bm{w}_{k}}}\mathrm{d}h$
$=P_{\mathrm{NA}}\frac{12\bm{w}_{k}^{\mathrm{T}}\mathbf{R}_{\bm{p}}\bm{w}_{k}}{\left(K+1\right)\left(K+2\right)}$
$\quad+\bm{w}_{k}^{\mathrm{T}}\mathbf{J}^{-1}\bm{w}_{k}\Gamma_{\frac{3}{2}}\left(\frac{\tilde{h}^{2}}{8\bm{w}_{k}^{\mathrm{T}}\mathbf{J}^{-1}\bm{w}_{k}}\right),$ (58)
where the approximate equality is due to the tight upper bound of the Q-function, $\tilde{h}_{k}$ represents the boundary between the asymptotic region and the non-asymptotic region, which is given by
$\tilde{h}_{k}^{2}\leq 2\bm{w}_{k}^{\mathrm{T}}\mathbf{J}^{-1}\bm{w}_{k}\sum_{i=0}^{P-1}{}\left(\frac{2N_{\mathrm{BS}}N\rho_{i}}{2+N_{\mathrm{BS}}N\rho_{i}}\right)^{2},$ (59)
with
$\tilde{h}_{k}^{2}=\min\left\{2\bm{w}_{k}^{\mathrm{T}}\mathbf{J}^{-1}\bm{w}_{k}\sum_{i=0}^{P-1}{}\left(\frac{2N_{\mathrm{BS}}N\rho_{i}}{2+N_{\mathrm{BS}}N\rho_{i}}\right)^{2},K\zeta_{k}^{2}\right\},$ (60)
and $\Gamma_{a}\left(\cdot\right)$ represemts the incomplete gamma function with parameter $a$.
As can be observed from (IV), the ZZB constitutes a global lower bound. In the regime of small $\mathbf{{\bm{\delta}}}$, the ZZB is dominated by the asymptotic CRB. As the SNR $\rho_{i}$ decreases, the non-asymptotic component $P_{\mathrm{NA}}$ of (IV) increasingly dominates and converges to the parameter covariance.
V Simulation Results
In this section, simulation results are demonstrated to examine the performance of the proposed channel estimation methods and the comparison with benchmarks. Table I gives the simulation parameters. The SNR is defined as $\mathrm{SNR}\triangleq\frac{\left\|\bm{\mathcal{Y}}\right\|_{\mathrm{F}}^{2}}{\left\|\bm{\mathcal{N}}\right\|_{\mathrm{F}}^{2}}$. The maximum MS velocity is set to 300 km/h, resulting in a maximum Doppler frequency of $\nu_{\max}=$ 4167 Hz. As the maximum Doppler tap is $\alpha_{\max}=\nu_{\max}T=0.14$, we set $c_{1}=\frac{1}{2M}$ to reduce the pilot overhead.
TABLE I: Simulation Parameters
Parameter Values
Carrier frequency (GHz) 15
Subcarrier spacing (kHz) 30
Total number of subcarriers and number of symbols for pilot $(M,N)$ (1024,11)
Length of CPP $M_{\text{CPP}}$ 72
Guard interval 72
Number of BS antennas 16
Number of MS antennas 1
The number of propagation paths 5
Maximum MS velocity (km/h) 300
Symbol Modulation Scheme 16/ 64/ 256 QAM
V-A Channel Parameters Estimate Performance
(a) The MSEs performance versus SNR.
Figure 6: The NMSE performance versus SNR.
Fig. 6(a) shows the MSEs performance versus SNR. The total number of subcarriers is set to $M=512$. The length of CPP is set to $M_{\text{CPP}}=45$. The rectified ALS [20] and the Newtonized orthogonal matching pursuit (NOMP) [28] are selected as benchmarks, which are denoted as “Rec-ALS” and “NOMP”, respectively. Specifically, the CRB and ZZB serve as benchmarks for AoA, Doppler, and delay estimation, while the LMMSE is used for channel gain. It can be observed that in the low SNR regime (SNR << 5 dB), the ZZBs for AoA and delay effectively capture the rapid degradation of estimation error. When SNR == -10 dB, the MSE of the proposed algorithm is comparable to the ZZB. However, in the SNR range of -5 dB to 5 dB, a noticeable gap exists between the MSE of the proposed algorithm and the ZZB. This discrepancy stems from the fact that the ZZB remains a performance bound derived from the ML criterion [29], and its behavior in the non-asymptotic regime is predominantly governed by the array gain $N_{\mathrm{BS}}N\rho_{i}$. While the proposed subspace-based algorithm for off-grid parameter estimation offers significantly higher computational efficiency, it generally yields suboptimal performance relative to ML estimation (MLE). Furthermore, the stochastic interaction between MSE for subspace-based method and noise is complicated and cannot be straightforwardly formulated as a function of array gain $N_{\mathrm{BS}}N\rho_{i}$. In addition, due to the relatively short symbol duration $T_{\mathrm{sym}}$, the MSE of Doppler estimation in the low-SNR regime ($\leq$5 dB) is considerably larger than both the ZZB and the CRB. While the SNR $\geq$ 10 dB, the ZZBs asymptotically converges to CRB. The MSEs of the proposed algorithm for AoA, Doppler and delay become close to ZZB and CRB. Finally, the MSE of the channel gain exhibits a similar trend with respect to the LMMSE: it remains relatively close in both the low SNR and medium-to-high SNR regions, while showing a larger gap in the low-to-medium SNR range.
V-B Communication Performance Comparison with Existing Methods
In this subsection, we examine the communication performance for the proposed channel estimation algorithms and the benchmarks. The NMSE performance is firstly examined. The CP-ALS [27], the rectified ALS [20] and the NOMP with perfect channel gain [28] are selected as benchmarks, which are denoted as “CP-ALS”, “Rec-ALS” and “NOMP-Oracle”, respectively. The NMSE is defined as
[…]
Fig. 6 illustrates the NMSE performance for the proposed methods and benchmarks. From Fig. 6, it can be observed that the proposed method achieves better accurate channel estimation performance than the benchmarks. The NMSE performance of the RecALS method is also relatively close to that of the proposed method. The main reason is both the proposed method and the RecALS method effectively exploit the rotation-invariant properties in both the spatial and time domains. Nevertheless, the inherently weak convergence of alternative LS process limits the average NMSE performance of RecALS. In addition, the NOMP algorithm based on [28] performs unsatisfactorily in multipath channel estimation. The reason is that the NOMP algorithm discards the off-diagonal blocks of the Hessian matrix during global iterations, which means the algorithm essentially still performs local iterations for each path separately. As a result, the iteration process may incurs significant bias.
(a) The BER performance versus SNR for different order QAM modulations.
The BER performance versus SNR for the proposed method and benchmarks for different order QAM modulations is demonstrated in Fig. 9(a). From Fig. 9(a), the BER performance of the proposed method outperforms that of the benchmarks among all the three QAM modulation orders. In addition, under high-order modulations of 64 QAM and 256 QAM, the proposed algorithm achieves a BER close to $10^{-2}$ and exceeding $10^{-3}$ at high SNRs, respectively, demonstrating the potential of the proposed method in MIMO-AFDM scenarios with high-order QAM modulation.
Figure 9: The SE performance versus SNR.
In Fig. 9, the SE performance versus SNR is shown. It can be seen that the SE performance of the proposed method is relatively close to that of the RecALS algorithm. In the SNR of 0-5 dB range, RecALS exhibits slightly better SE performance. However, when the SNR exceeds 5 dB, the proposed method achieves higher SE.
Figure 10: The average CPU running time versus SNR.
To evaluate the computational efficiency of different AFDM channel estimation algorithms, the average CPU running time versus SNR is demonstrated in Fig. 10. As shown in Fig. 10, the NOMP algorithm requires relatively high CPU time, as it needs to update the gradient matrix, the Hessian matrix, and the objective function at each iteration, all of which contribute to its high computational complexity. Furthermore, compared with NOMP, both the classical CP-ALS algorithm and the RecALS algorithm reduce the computational time by approximately one order of magnitude. This highlights the advantage of tensor decomposition-based methods in achieving low complexity when processing high-dimensional signals. Finally, since the proposed algorithm is based on Vandermonded structure $T$-decomposition and involves no iterative process, the computational time of the proposed algorithm is further reduced by approximately one order of magnitude compared with CP-ALS and RecALS algorithm. The low computational complexity of the proposed algorithm makes it highly suitable for deployment in high-mobility scenarios.
Figure 11: The NMSE performance versus the number of propagation paths, the SNR is 20 dB.
To further investigate the estimation performance of the algorithms under channels with different propagation scenarios, Fig. 11 shows the NMSE performance versus the number of propagation paths, where the SNR is set to 20 dB. It can be observed from Fig. 11 that the NMSE performance of the classical CP-ALS algorithm is mostly affected with the increase in the number of channel propagation paths. As the path number exceeds four, the NMSE performance of the CP-ALS deteriorates sharply. The main reason is the poor convergence behavior of thhe CP-ALS when processing the high-rank tensors [23]. Furthermore, the performance degradation of the RecALS algorithm is less pronounced as the number of paths increases. This is attributed to the fact that, unlike the classical CP-ALS, the RecALS fully leverages the Vandermonde factor matrices within the tensor. Finally, the proposed algorithm incurs less performance loss compared to the CP-ALS and the RecALS. In contrast to these two ALS-based algorithms, the proposed $T$-decomposition-based algorithm demonstrates superior robustness in processing high-rank tensor signals.
Figure 12: The BER performance versus the number of propagation paths, the SNR is 20 dB.
Fig. 12 evaluates the BER performance versus the number of propagation paths. As illustrated in Fig. 12, only the proposed algorithm achieves a reduction in BER as the number of propagation paths increases. As mentioned earlier, the proposed algorithm exhibits the least NMSE degradation with the increase in paths. This characteristic secures a robust diversity gain, which directly results in an improved BER performance.
V-C Versus OTFS and OFDM
To further investigate the performance comparison between AFDM, OTFS and OFDM in more realistic environments, we simulate their BER and SE performance under conditions of fractional delay and Doppler. The maximum velocity is set to 500 km/h, the total number of subcarriers is set to $M=512$, and the number of symbols in the OTFS frame is $N=13$, resulting in a maximum Doppler tap of $3.43$. According to [36], the guard interval is equal to the ceiling of the maximum Doppler tap. This implies that the pilot length in the time domain extends to NN (i.e., the full frame duration). The number of pilots used for AFDM is set to $N_{p}=7$  [^1] For the proposed algorithm, it is only necessary when the number of pilot symbols satisfies $N_{p}\geq P$.. The other parameters are the same as in Table I.
Figure 13: The BER performance versus SNR.
Fig. 13 demonstrates the BER performance versus the SNR for AFDM, OTFS and OFDM. It can be observed from the simulation results that, under identical time-frequency resource configurations, the BER performance of OTFS is marginally superior to that of AFDM. This is attributed to the fact that OTFS employs 2D inverse symplectic finite Fourier transform (ISFFT) modulation. Compared to AFDM, which employs 1D DAFT modulation, the impulse response of OTFS in the 2D delay-Doppler domain exhibits higher energy concentration. Consequently, OTFS demonstrates enhanced robustness against noise, leading to better BER performance.
Figure 14: The SE performance versus SNR.
Fig. 14 shows the SE performance versus SNR for AFDM, OTFS and OFDM. The SE is calculated as
$\mathrm{SE}=(1-\gamma)\log_{2}\left(1+\mathrm{SINR}\right),$ (62)
where $\gamma$ denotes the pilot overhead in time-frequency resources. We have $\gamma=\gamma^{\mathrm{A}}$ for AFDM and $\gamma=\gamma^{\mathrm{O}}$ for OTFS, and
$\gamma^{\mathrm{A}}$ $=2N_{p}M_{g}/\left(NM\right),$ (63)
$\gamma^{\mathrm{O}}$ $=2M_{g}/M,$ (64)
[…]
VI Conclusion
In this paper, we investigated the channel estimation for MIMO-AFDM systems and analyzed the input-output relationship for MIMO-AFDM systems with the consideration of fractional delay and Doppler. To tackle with the estimation of fractional delay and Doppler, we proposed a T-AF domain embedded pilot structure. By leveraging the magnitude of AF domain and the rotational invariance property of time domain, we proposed an efficient Vandermonde structure $T$ decomposition-based channel estimation algorithm. Compared to the existing parameter algorithms, the proposed algorithm significantly reduce the computational complexity. To evaluate the MSE performance, we derived Ziv-Zakai bound for the 3D MIMO-AFDM signal model. Simulation results showed that the proposed algorithm significantly outperforms the benchmarks at medium-to-high SNR.
Appendix A Proof of Theorem 1
To physically unveil the signal interaction in doubly dispersive channels, we adopt a derivation approach for the ambiguity function similar to [38] and [39]. Although the derivation started in an integral form, the results take into account the process of discrete sampling and the application of the DAFT. As shown in the formula in the text, the derivation results of this paper are in complete agreement with the signal model in [8] under integer delay scenarios.
Based on (2)-(7), we have the following STAF domain symbols $Y\left[n_{\mathrm{BS}},n,m\right]$ as (65), which is shown at the bottom of this page. Then, the STAF domain signal $Y\left[n_{\mathrm{BS}},n,m\right]$ can be obtained via discrete sampling the ambiguity function as (66), which is shown at the bottom of the next page.
$A_{g_{\mathrm{rx}},r_{n_{\mathrm{BS}}}}\left(t,f\right)\left|{}_{t=nT_{\mathrm{sym}},f=m\Delta f}\right$. (65)
$\quad=\int_{nT_{\mathrm{sym}}+T_{\mathrm{CPP}}}^{nT_{\mathrm{sym}}+T_{\mathrm{sym}}}{}e^{-j\left(\frac{d}{2b}\left(t^{\prime}-t-T_{\mathrm{CPP}}\right)^{2}+2\pi\left(m\Delta f\right)\left(t^{\prime}-t-T_{\mathrm{CPP}}\right)+\frac{a}{2b}\left(m\Delta f\right)^{2}\right)}g_{\mathrm{rx}}^{\ast}\left(t^{\prime}-t\right)r_{n_{\mathrm{BS}}}\left(t^{\prime}\right)\mathrm{d}t^{\prime}$
$\quad=\sum_{n^{\prime}=0}^{N-1}{}\sum_{m^{\prime}=0}^{M-1}{}\int_{nT_{\mathrm{sym}}+T_{\mathrm{CPP}}}^{nT_{\mathrm{sym}}+T_{\mathrm{sym}}}{}e^{-j\left(\frac{d}{2b}\left(t^{\prime}-t-T_{\mathrm{CPP}}\right)^{2}+2\pi\left(m\Delta f\right)\left(t^{\prime}-t-T_{\mathrm{CPP}}\right)+\frac{a}{2b}\left(m\Delta f\right)^{2}\right)}g_{\mathrm{rx}}^{\ast}\left(t^{\prime}-t\right)$
$\quad\quad\times\left[\iiint{}\left[\mathbf{a}_{N_{\mathrm{BS}}}\left(\theta\right)\right]_{n_{\mathrm{BS}}}h\left(\tau,\nu,\theta\right)e^{j2\pi\nu\left(t^{\prime}-\tau\right)}x_{n^{\prime},m^{\prime}}\phi_{m^{\prime}}\left(t^{\prime}-n^{\prime}T_{\mathrm{sym}}-\tau\right)g_{\mathrm{tx}}\left(t^{\prime}-n^{\prime}T_{\mathrm{sym}}-\tau\right)\mathrm{d}\tau\mathrm{d}\nu\mathrm{d}\theta\right]\mathrm{d}t^{\prime}$
$\quad=\sum_{m^{\prime}=0}^{M-1}{}x_{n,m^{\prime}}\iiint{}\left[\mathbf{a}_{N_{\mathrm{BS}}}\left(\theta\right)\right]_{n_{\mathrm{BS}}}h\left(\tau,\nu,\theta\right)$
$\quad\quad\times\int_{nT_{\mathrm{sym}}+T_{\mathrm{CPP}}}^{nT_{\mathrm{sym}}+T_{\mathrm{sym}}}{}e^{j2\pi\nu\left(t^{\prime}-\tau\right)}e^{-j\left(\frac{d}{2b}\left(t^{\prime}-t-T_{\mathrm{CPP}}\right)^{2}+2\pi\left(m\Delta f\right)\left(t^{\prime}-t-T_{\mathrm{CPP}}\right)+\frac{a}{2b}\left(m\Delta f\right)^{2}\right)}\phi_{m^{\prime}}\left(t^{\prime}-nT_{\mathrm{sym}}-\tau\right)\mathrm{d}t^{\prime}$
$\quad=\sum_{m^{\prime}=0}^{M-1}{}x_{n,m^{\prime}}\sum_{i=1}^{P}{}\alpha_{i}\left[\mathbf{a}_{N_{\mathrm{BS}}}\left(\theta\right)\right]_{n_{\mathrm{BS}}}$
$\quad\quad\times\int_{nT_{\mathrm{sym}}+T_{\mathrm{CPP}}}^{nT_{\mathrm{sym}}+T_{\mathrm{sym}}}{}e^{j2\pi\nu_{i}\left(t^{\prime}-\tau_{i}\right)}e^{-j\left(\frac{d}{2b}\left(t^{\prime}-t-T_{\mathrm{CPP}}\right)^{2}+2\pi\left(m\Delta f\right)\left(t^{\prime}-t-T_{\mathrm{CPP}}\right)+\frac{a}{2b}\left(m\Delta f\right)^{2}\right)}\phi_{m^{\prime}}\left(t^{\prime}-nT_{\mathrm{sym}}-\tau_{i}\right)\mathrm{d}t^{\prime}$
$Y\left[n_{\mathrm{BS}},n,m\right]$ $=\sum_{m^{\prime}=0}^{M-1}{}x_{n,m^{\prime}}\sum_{i=1}^{P}{}\tilde{\alpha}_{i}\left[\mathbf{a}_{N_{\mathrm{BS}}}\left(\theta\right)\right]_{n_{\mathrm{BS}}}\frac{1}{M}\sum_{p=M_{\mathrm{CPP}}}^{M+M_{\mathrm{CPP}}-1}{}e^{j2\pi\nu_{i}\frac{T}{M}p}e^{-j2\pi c_{1}\left(p-M_{\mathrm{CPP}}\right)^{2}}e^{-j2\pi\frac{m\left(p-M_{\mathrm{CPP}}\right)}{M}}e^{-j2\pi c_{2}m^{2}}$ (66)
$\quad\times e^{j2\pi c_{1}\left(p-l_{i}-\iota_{i}-M_{\mathrm{CPP}}\right)^{2}}e^{j2\pi\frac{m^{\prime}\left(p-l_{i}-\iota_{i}-M_{\mathrm{CPP}}\right)}{M}}e^{j2\pi c_{2}{m^{\prime}}^{2}}$
$=\sum_{m^{\prime}=0}^{M-1}{}x_{n,m^{\prime}}\sum_{i=1}^{P}{}\tilde{\alpha}_{i}\left[\mathbf{a}_{N_{\mathrm{BS}}}\left(\theta\right)\right]_{n_{\mathrm{BS}}}e^{j2\pi\nu_{i}nT_{\mathrm{sym}}}H_{i}\left[m,m^{\prime}\right]$
Appendix B Proof of Theorem 2
Using equation $\frac{\partial\mathbf{X}\left(s\right)^{-1}}{\partial s}=-\mathbf{X}\left(s\right)^{-1}\frac{\partial\mathbf{X}\left(s\right)}{\partial s}\mathbf{X}\left(s\right)^{-1}$, the second-order partial derivative of $\mu\left(s;\mathbf{{\bm{\delta}}}\right)$ w.r.t. $s$ is given by
[…]
$\quad\left.\left.\times\left(\mathbf{R}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}+\mathbf{{\bm{\delta}}}}^{-1}-\mathbf{R}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}}^{-1}\right)\right)^{2}\right).$ (67)
We firstly examine the values of $\mu\left(s;\mathbf{{\bm{\delta}}}\right)$ and $\frac{\partial^{2}\mu\left(p;\mathbf{{\bm{\delta}}}\right)}{\partial s^{2}}$ in the asymptotic regime, i.e., in the vicinity of $\mathbf{{\bm{\delta}}}=\mathbf{0}_{3P}$. Performing the first-order taylor expansion of $\mathbf{R}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}+\mathbf{{\bm{\delta}}}}$ as
$\mathbf{R}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}+\mathbf{{\bm{\delta}}}}\approx\mathbf{R}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}}+\underset{\Delta\mathbf{R}}{\underbrace{\sum_{i}{}{\bm{\delta}}_{i}\frac{\partial\mathbf{R}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}}}{\partial p_{i}}\;\mid_{\mathbf{{\bm{\varphi}}}=\bm{p}}}}.$ (68)
Then, we can expand the function $\mu\left(s;\mathbf{{\bm{\delta}}}\right)$ by applying the approximation $\left(\mathbf{I}+\mathbf{X}\right)^{-1}\approx\mathbf{I}-\mathbf{X}$ and $\ln\left|\mathbf{I}+\mathbf{X}\right|\approx\mathrm{Tr}\left(\mathbf{X}\right)-\frac{1}{2}\mathrm{Tr}\left(\mathbf{X}^{2}\right)$ as
$\mu\left(s;\mathbf{{\bm{\delta}}}\right)$ $\approx-\ln\left|s\left(\mathbf{I}+\mathbf{R}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}}^{-1}\Delta\mathbf{R}\right)^{-1}+\left(1-s\right)\mathbf{I}\right|$
$\quad-s\ln\left|\mathbf{I}+\mathbf{R}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}}^{-1}\Delta\mathbf{R}\right|$
$\approx s\mathrm{Tr}\left(\mathbf{R}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}}^{-1}\Delta\mathbf{R}\right)+\left(\frac{s^{2}}{2}-s\right)\mathrm{Tr}\left(\left(\mathbf{R}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}}^{-1}\Delta\mathbf{R}\right)^{2}\right)$
$\quad-s\mathrm{Tr}\left(\mathbf{R}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}}^{-1}\Delta\mathbf{R}\right)+\frac{s}{2}\mathrm{Tr}\left(\left(\mathbf{R}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}}^{-1}\Delta\mathbf{R}\right)^{2}\right)$
$=\frac{s^{2}-s}{2}\mathrm{Tr}\left(\left(\mathbf{R}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}}^{-1}\Delta\mathbf{R}\right)^{2}\right)=\frac{s^{2}-s}{2}\mathbf{{\bm{\delta}}}^{\mathrm{T}}\mathbf{J{\bm{\delta}}}.$ (69)
Hence, we have $\left.\mu\left(s;\mathbf{{\bm{\delta}}}\right)\right|_{p=\frac{1}{2}}=-\frac{1}{8}\mathbf{{\bm{\delta}}}^{\mathrm{T}}\mathbf{J{\bm{\delta}}}$. Similarly, $\frac{\partial^{2}\mu\left(s;\mathbf{{\bm{\delta}}}\right)}{\partial s^{2}}$ can be rewritten as
[…]
$\quad\times\left.\left.\left(\left(\mathbf{R}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}}+\Delta\mathbf{R}\right)^{-1}-\mathbf{R}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}}^{-1}\right)\right)^{2}\right)$
$\approx\mathrm{Tr}\left(\left(\left[\mathbf{R}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}}+\Delta\mathbf{R}\right]\mathbf{R}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}}^{-1}\Delta\mathbf{RR}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}}^{-1}\right)^{2}\right)$
$\approx\mathrm{Tr}\left(\left(\Delta\mathbf{RR}_{\mathbf{y}|\mathbf{{\bm{\varphi}}}}^{-1}\right)^{2}\right)$
$=\mathbf{{\bm{\delta}}}^{\mathrm{T}}\mathbf{J{\bm{\delta}}}.$ (70)
Next, we analyze the behavior of the function $\mu\left(s;\mathbf{{\bm{\delta}}}\right)$ and $\frac{\partial^{2}\mu\left(s;\mathbf{{\bm{\delta}}}\right)}{\partial s^{2}}$ in the non-asymptotic regime. When $\mathbf{{\bm{\delta}}}$ grows large, the array steering vector $\mathbf{a}_{N_{\mathrm{BS}}}\left(\theta_{i}\right)$ and $\mathbf{a}_{N_{\mathrm{BS}}}\left(\theta_{i}+{\bm{\delta}}_{\theta_{i}}\right)$ exhibit asymptotic orthogonality [33]. Similar trend also appears for $\mathbf{b}_{N}\left(\nu_{i}\right)$ and $\mathbf{b}_{N}\left(\nu_{i}+{\bm{\delta}}_{\nu_{i}}\right)$. Furthermore, as $\mathbf{{\bm{\delta}}}_{\eta_{i}}$ increases, the main-lobe positions of $\mathbf{c}_{m_{\mathrm{pilot}}}\left(\nu_{i},\eta_{i}\right)$ and $\mathbf{c}_{m_{\mathrm{pilot}}}\left(\nu_{i},\eta_{i}+\mathbf{{\bm{\delta}}}_{\eta_{i}}\right)$ become misaligned, and the correlation $\left|\mathbf{c}_{m_{\mathrm{pilot}}}\left(\nu_{i},\eta_{i}\right)^{\mathrm{H}}\mathbf{c}_{m_{\mathrm{pilot}}}\left(\nu_{i},\eta_{i}+\mathbf{{\bm{\delta}}}_{\eta_{i}}\right)\right|$ gradually approaches zero. Therefore, it can be inferred that, as $\mathbf{{\bm{\delta}}}$ increases asymptotically, the subspaces spanned by respectively become asymptotically orthogonal. The two covariance matrices contribute equally to the eigenvalues in the signal subspace, both being $\frac{1}{\left(\sigma_{n}^{2}+N_{\mathrm{BS}}N\sigma_{i,\mathrm{PL}}^{2}\right)}$, while the eigenvalue contributions in the remaining noise subspace are $\frac{1}{\sigma_{n}^{2}}$. Thus, we have
$\left.\mu\left(s;\mathbf{{\bm{\delta}}}\right)\right|_{s=\frac{1}{2}}\approx\sum_{k}{}\ln\left[\frac{4\left(1+N_{\mathrm{BS}}N\rho_{i}\right)}{\left(2+N_{\mathrm{BS}}N\rho_{i}\right)^{2}}\right],$ (71)
and
$\left.\frac{\partial^{2}\mu\left(s;\mathbf{{\bm{\delta}}}\right)}{\partial s^{2}}\right|_{s=\frac{1}{2}}=2\sum_{i=0}^{P-1}{}\left(\frac{2N_{\mathrm{BS}}N\rho_{i}}{2+N_{\mathrm{BS}}N\rho_{i}}\right)^{2}.$ (72)
Substituting (69)-(72) into (IV), we obtain (50), which completes the proof.